OpenStudy (anonymous):

-6p+7=3(2p-3)-4(-10+4p)

OpenStudy (anonymous):

First simplify and distribute: -6p+7=3(2p-3)-4(-10+4p) -6p + 7 = 6p - 9 + 40 - 16p Combine like terms on both sides: -6p + 7 = -10p + 31 Add 10p to both sides -6p + 10p +7 = -10p+10p+31 Simlify 4p + 7 = 31 Subtract 7 from both sides 4p = 24 Divide p=6

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Yep. :)